General Meeting Summary 5/10/17
Jump to navigation
Jump to search
DETECTORS AT JLAB (Arthur)
- SHMS Aerogel detector - installed and ready, waiting for aerogel tray exchange for SP-11 tests
- Tray exchange tentatively scheduled for end of May - techs are currently busy with the NPS/GEM clean room construction
- Spare PMTs have been relocated to Physics Storage (Picture of the storage arrangements)
- HMS Aerogel detector - discussion about total PMT failure
- Based on testing over the last week, it was found that all 14 PMTs are dead
- Possible causes of this failure:
- Helium leak in the hall or spectrometer enclosure - this would be of concern also for the SHMS aerogel detector PMTs
- PMT intrinsic failure, e.g. lost vacuum over time (the HMS aerogel detector was installed in 2003)
- Next steps:
- Investigate the HMS aerogel detector PMTs in more detail - check out photocathode etc.
- Hall C plans to install a helium leak detection system - follow up on that and any other information
- Neutral Particle Spectrometer (NPS) - construction progressing, need information on mechanical space required for signal/HV connectors for decision on HV divider design
- Action item: Arthur will take a picture of the back of the 3x3 NPS prototype to illustrate the mechanical space requirements
DETECTORS AT CERN (Rachel)
- CMS/CERN has first beams! - saw first cosmic muon tracks through the detectors
- Pixel detectors are on only during stable beams - expected around 25 May. Taking opportunistic data at the moment for diagnostics
- one of the current tasks is the alignment of the detector. Alignment measurements were done previously with cosmics and magnet off. Now have to redo with magnet on as the detector might have moved
- Physics with CMS could start in early June
PBWO4 STUDIES AT CUA (Richard, Andres)
- Discussion about crystal prototype
- components for final mechanical calibration have been constructed. Expect mechanical calibrations to be done within ~1week
- APD readout - discussion about managing background noise
- Arthur can temporarily provide a pulse generator that should be more stable than what Andres is using at the moment
- Discussion that even with perfect pulse generator, LED may still not be stable (depends on type, voltage etc.) - see Hamlet's recent studies of HV dividers with LEDs. It may be ok at the 10% level. May need to think about a setup including a reference PMT if want accuracy of ~1%.
EXPERIMENT PROPOSALS AND DATA ANALYSIS
- TCS - decided to postpone submission until next year
- TDIS Kaon - run group proposal in preparation for 12 GeV JLab, ongoing studies for measurements at an EIC to be presented at Argonne National Lab June 1-2 (http://www.phy.anl.gov/theory/pieic2017/index.html)
- DVCS-3 - deadtime analysis of recently acquired data ongoing
NEXT MEETING: WEDNESDAY 17 MAY at 10:00 AM (EDT)